# Yumeng Wang2    Zhiyuan Fan2    Qingyun Wang1    May Fung2  , Heng Ji1∗ 1

### v1

@@ 5 Conclusion
Table 4: We investigate the cross-dataset generalizability. The table shows the result of training on MEDQA and testing on X-CSQA, or training on X-CSQA and testing on MEDQA.
@@ Appendix A Training and inference configuration
We set $m=3$ when sampling responses for each of the base models. We finally obtained 17244 and 2168 preference pairs from MEDQA and X-CSQA datasets, respectively. We used LoRA Hu et al. (2021) Fine-tuning method for DPO and SFT training. The training parameters are listed in Table 5. The inference parameters are shown in Table 6. All the experiments are performed on NVIDIA A100-SXM-80GB GPUs.
@@
MEDQA(%) X-CSQA(%)
Model EN CN FR EN CN FR
@@
Table 8: Percentage of each language in the final training data.
MEDQA US MEDQA CN-ZH
Model Native EN EN2CN EN2FR AVG Native CN CN2EN CN2FR AVG
@@
Appendix B Detailed use of the training dataset
B.1 Data source
B.2 Statistics of the training datasets
Table 7 shows the percentages of positive samples for each language across task settings. English indeed tasks up the largest portion of the positive samples, but there are still considerable amounts of Chinese and French samples.
B.3 Full result of MEDQA dataset

### v2

@@ 4.2 SFT and DPO with ground truth
Table 3: We investigate the cross-dataset generalizability. The table shows the result of training on MEDQA and testing on X-CSQA, or training on X-CSQA and testing on MEDQA. Both settings surpass the baseline.
4.3 Generalizability
Cross-dataset generalizability
To evaluate the generalizability, we conduct cross-dataset experiments by training models on X-CSQA and testing them on MEDQA, and vice versa. Table 3 reveals that while the out-of-domain accuracy falls below the in-domain accuracy, it consistently exceeds the in-domain performance of the SFT baseline. This underscores the capability of CALM-trained models to provide multilingually consistent answers, even when faced with unseen tasks or domains. These findings suggest that CALM enhances in-domain performance and fosters robustness across different types of domains.
Cross-lingual generalizability
We implement CALM training sequentially, beginning with English and incrementally adding French and Chinese, progressing from high-resource to low-resource languages. At each step, we evaluate test accuracy across all languages. To assess CALM’s effectiveness in untrained languages, we include Japanese, Italian, and German in the test set, none of which were included during training. In Table 10 in the Appendix, CALM demonstrates greater effectiveness as more languages participate in majority voting. Notably, even untrained languages exhibit accuracy improvements, suggesting that CALM’s alignment mechanism fosters a unified understanding of knowledge across languages, thereby enhancing overall comprehension. This aligns with She et al. (2024), which similarly observe cross-lingual generalizability in multilingual reasoning tasks.
5 Conclusion
We introduce CALM, a novel framework to facilitate the alignment of LLM’s knowledge across different languages. We observe that CALM is more effective when more languages are involved in the training, due to internal knowledge aggregation. Additionally, CALM outperforms ground truth DPO and SFT. It shows that although some of the positive samples are factually incorrect, they also contribute to the accuracy gain in CALM, possibly because more consistent answers often have better task understanding and can lead the model towards more correct answers. Through comprehensive experiments, we demonstrate the effectiveness of CALM in achieving robust cross-lingual knowledge alignment.
@@ Appendix A Training and inference configuration
We set $m=3$ when sampling responses for each of the base models. We finally obtained 17244 and 2168 preference pairs from MEDQA and X-CSQA datasets, respectively. We used LoRA Hu et al. (2021) Fine-tuning method for DPO and SFT training. The training parameters are listed in Table 5. The inference parameters are shown in Table 6. All the experiments are performed on NVIDIA A100-SXM-80GB GPUs. We utilize the Llama3-8B-Instruct and Mistral-7B-Instruct model from LlamaFactory Zheng et al. (2024) framework for training and testing.
@@
Model MEDQA(%) X-CSQA(%)
EN CN FR EN CN FR
@@
Table 8: Percentage of Chinese, French and English language in final CALM training data.
Model MEDQA US MEDQA CN-ZH
Native EN EN2CN EN2FR AVG Native CN CN2EN CN2FR AVG
@@
Appendix B Detailed use of the training dataset
B.1 Data source
B.2 Statistics of the training datasets
Table 7 and Table 8 shows the percentages of positive samples for each language across task settings. English indeed tasks up the largest portion of the positive samples, but there are still considerable amounts of Chinese and French samples.
B.3 Full result of MEDQA dataset